A balloon of mass M is descending at a constant acceleration Alpha when a mass M is released from the balloon it starts rising with the same acceleration alpha assuming that its volume does not change What is the value of m

Grade:11

1 Answers

Shivam Shekhar
6 years ago
Let F be the pulling force ....Given, accn.=a( I am taking it instead alpha)..given that mass of the baloon is M and the mass thrown out m...Through FBD when it goes dow with accn. a ,then..Mg-F=Ma---iAnd when it goes upwards then the accn.is also a;soF-(M-m)g=(M-m)a----iiSolve the equations and the answer ism=2Ma/(g+a)